# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v1

@@ 3.2 Base Neural Architecture for UV-Parameterization Learning
To emulate the sequence of operations an artist would typically perform when constructing a UV atlas for a 3D mesh, such as stretching a canvas, placing it over the 3D surface, adding surface cuts where needed, and finally flattening it back to the plane, we build upon the idea of a deforming a 2D grid Yang et al. (2018) by leveraging bi-directional cycle mapping backbone, proposed in Zhang et al. (2024a; b); Srinivasan et al. (2024); Zhao et al. (2025). Intuitively, the bi-directional cycle mapping enforces that projecting a 3D surface patch into 2D UV coordinates and re-projecting those UVs back onto the 3D mesh yields the original data. Therefore, during the learning process, any seams, local stretches, distortions, or semantic misalignments show up as reconstruction errors. Then, by minimizing these errors, the model is guided to produce an approximately invertible and coherent parameterization that minimizes distortion. To build such a mapping, inspired by recent advances in global parameterization networks Zhao et al. (2025); Srinivasan et al. (2024), our backbone is composed of four geometrically interpretable subnetworks: (i) deforming network, which starts from a regular 2D grid and adaptively deforms it into candidate UV coordinates; (ii) wrapping network, which maps those 2D UV candidates to the 3D surface by smoothly bending the deformed grid to conform to the input mesh, bridging the 2D and 3D domains; (iii) cutting network, which operates on the wrapped 3D surface to open the closed geometry by placing seams that reduce distortion for subsequent flattening; and (iv) unwrapping network, which projects the resulting open 3D surface back to 2D while preserving smoothness and geometric coherence. In this design, Each subnetwork is a pointwise MLP applied to mesh vertices, and the whole architecture is trained end-to-end in an unsupervised fashion. The four subnetworks form two complementary branches enforcing 2D–3D–2D and 3D–2D–3D cycles for ensuring consistency across 2D and 3D domains. Coupled with differentiable losses that encourage bijectivity and minimize distortion, this backbone produces candidate UV parameterizations for the next stages. Further details on the loss design and the interaction of the four sub-networks within the two-branch bi-directional cycle are given in Appendix A.2.
@@ 3.4 Visibility-Aware UV Parameterization
Next, we define a threshold $\tau$ based on the UV domain size. Specifically, we compute the bounding square of the UV map with side length $L(Q)$, and set $\tau=\tau_{\text{scale}}\cdot L(Q)$ where $\tau_{\text{scale}}$ is a constant (e.g., $0.1$ by default). However, this maximization is not inherently differentiable, so we introduce mathematical modifications to ensure differentiability during the learning process. Specifically, instead of a hard decision, we assign each vertex a differentiable soft seam membership score: $s_{i}=\sigma\big(\beta\,(\eta_{i}-\tau)\big)$ where $\sigma(\cdot)$ is the sigmoid function and $\beta$ controls the sharpness of the transition. The maximization problem can then be formulated as a differentiable function:
@@
Visibility-Aware Seam Loss. After identifying soft seam memberships $s_{i}\in[0,1]$ for each vertex $v_{i}$, we encourage seams to lie in less visible regions of the surface. As a visibility proxy, we use per-vertex ambient occlusion values $AO_{i}\in[0,1]$, computed in the first step, where 0 indicates fully visible and $1$ indicates fully occluded. The loss is defined as the weighted average of occlusion values, using soft seam memberships as weights:
@@
where $\varepsilon$ is a small constant to avoid division by zero. Intuitively, if the network assigns high seam weight $s_{i}$ to highly visible vertices (low $AO_{i}$), the loss is large. Conversely, assigning seam weight to occluded vertices (high $AO_{i}$) reduces the loss. Therefore, minimizing $\mathcal{L}_{\text{AO}}$ guides seam placement toward regions of low exposure.
@@ 4 Experiments and Results
We evaluate our unsupervised semantic- and visibility-aware UV parameterization pipelines on a diverse collection of meshes and compare against several baselines. We provide both qualitative and quantitative comparisons to: FlexPara Zhao et al. (2025) (recent state-of-the-art global parameterization network), Autodesk Maya Autodesk, Inc. (2025) (industry-grade commercial software), Blender Blender Foundation (2025) (popular open-source modeling tool), and xatlas jpcy (2025) (widely-used C++ library for generating unique texture coordinates suitable for lightmap baking and texture painting). Figs. 1,  3,  4, and  5 show representative qualitative results, and Tables. 1 and  2 presents quantitative results of our semantic-aware and visibility-aware pipelines against baselines (additional results are provided in Appendix A.3). In the following, we detail each experiment and its corresponding results.
Comparative Evaluation of Semantic-Aware UV Parameterization. To evaluate the performance of the proposed semantic-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025), Autodesk Maya’s Automatic UV generator Autodesk, Inc. (2025), Blender’s SmartUV Blender Foundation (2025), and xatlas jpcy (2025). Fig. 3 shows a representative comparison on a single Bunny mesh; additional results on more diverse and complex meshes are provided in Figs. 11 and 12 in Appendix A.3.
@@
In Fig. 3 and Fig. 11, each row corresponds to one method. Semantic parts are color-coded, and columns show the rendered 3D mesh from multiple camera viewpoints. The rightmost column depicts the unwrapped UV coordinates, where UV islands are color-coded consistently with their corresponding semantic 3D parts in the rendered views. As shown, our method successfully identifies semantic parts of the input mesh, parameterizes them, and maps them into UV space with low distortion. This leads to cleaner, more interpretable charts that directly benefit downstream tasks such as texture painting and transfer. By contrast, the baseline methods fail to consistently capture and parameterize semantic parts of the 3D mesh.
@@
Comparative Evaluation of Visibility-Aware UV Parameterization. To evaluate the performance of the proposed visibility-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025) and OptCuts Li et al. (2018). Figs. 4 and  8 present representative comparisons on different meshes. Additional examples appear in Figs. 5,  9, and 10 in Appendix A.3. In Fig. 4 the top row visualizes per-vertex ambient occlusion (AO): lighter tones (e.g., yellow) indicate high exposure (more visible) while darker tones (e.g., purple) indicate low exposure (more occluded). A desirable parameterization therefore places cutting seams on darker, less-visible regions to reduce perceptible seam artifacts after texturing and rendering. The rows beneath show UV parameterizations produced by our method, FlexPara, and OptCuts (in that order). Qualitatively, our method allocates a larger fraction of seam geometry to low-exposure regions compared to the baselines. Table 1 (and Fig. 7 in Appendix A.3) quantitatively supports this observation: the report the mean AO values of seam vertices for each method, where lower scores indicate seams located in less-visible (more occluded) areas. As shown, our method consistently attains lower mean seam-AO, confirming that the visibility-aware objective successfully biases seam placement toward less-exposed surface regions. Also, Fig. 5 illustrates checkerboard texturing on different mesh objects using parameterizations from our method and from baselines. As shown, our method produces noticeably more seamless texture patterns when rendered from visible viewpoints, whereas FlexPara and OptCuts exhibit prominent cutting seams (highlighted in the zoomed-in insets of Fig. 5).
@@
Quantitative Comparison. We quantitatively evaluate both the semantic-aware and visibility-aware components of our pipeline on a collection of representative meshes. For the semantic-aware pipeline we report (i) Hamming distance (lower is better), the fraction of vertices whose labels disagree after optimal label matching to the reference (we use SAMesh Tang et al. (2024) outputs as a ground-truth); and (ii) Rand Index (RI) (higher is better), a pairwise agreement measure that counts how often vertex pairs are grouped or separated identically between the two labelings. For the visibility-aware pipeline we report Mean seam AO (lower is better in our AO convention, where 0 = occluded and 1 = exposed), the average ambient-occlusion value over seam vertices, which indicates how well seams are placed in occluded regions. In addition to these task-specific metrics, we evaluate geometric quality for all methods using conformality (angle-preservation) and equiareality (area-preservation) (higher is better) to ensure the parameterizations remain faithful to the original surface. Tables 1 and  2 summarize these metrics and compare our method against the baselines described in Sec. 4. Further discussion of the results in these two tables are provided in Appendix A.3.
@@ 5 Conclusions
Acknowledgments
This research was supported by Autodesk Research and the Autodesk AI Lab. We are especially thankful for the collaborative environment, technical resources, and insightful discussions that contributed to shaping this work.
@@ A.3 Additional Experiments and Results
We show expanded results, ablation studies, and discussions from the main paper featuring more surface parameterization and more viewpoints of different 3D objects in Figs. 714, and Tables. 4 and  5.
Additional Ablation Studies. In addition to the ablations reported in the main text, we further ablate the full framework by training identical models with and without the semantic and/or visibility objectives. Results in Table 5 show that adding semantic and/or visibility objectives to the base geometry-preserving UV parameterization model (Sec. 3.2 and Appendix A.2) preserves the model’s geometry-preserving properties: any loss in conformality or area preservation is minor. Also, these small quantitative degradations do not harm perceptual quality. Specifically, qualitative comparisons throughout the paper show the textured results remain competitive with baselines, despite slight drops in some numerical metrics.
@@
Table 4: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) while keeping the remainder of the partition-and-parameterize pipeline identical. Quantitatively, using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh.

### v2

@@ 3.2 Base Neural Architecture for UV-Parameterization Learning
To emulate the sequence of operations an artist would typically perform when constructing a UV atlas for a 3D mesh, such as stretching a canvas, placing it over the 3D surface, adding surface cuts where needed, and finally flattening it back to the plane, we build upon the idea of a deforming a 2D grid Yang et al. (2018) by leveraging bi-directional cycle mapping backbone, proposed in Zhang et al. (2024a; b); Srinivasan et al. (2024); Zhao et al. (2025). We emphasize that this stage adopts an existing previously proposed low-distortion parameterization backbone. We do not claim novelty for the backbone itself. Instead, our primary contribution lies in the design and integration semantic-aware and visibility-aware objectives as downstream tasks built on top of this established geometric foundation. Intuitively, the bi-directional cycle mapping enforces that projecting a 3D surface patch into 2D UV coordinates and re-projecting those UVs back onto the 3D mesh yields the original data. Therefore, during the learning process, any seams, local stretches, distortions, or semantic misalignments show up as reconstruction errors. Then, by minimizing these errors, the model is guided to produce an approximately invertible and coherent parameterization that minimizes distortion. To build such a mapping, inspired by recent advances in global parameterization networks Zhao et al. (2025); Srinivasan et al. (2024), our backbone is composed of four geometrically interpretable subnetworks: (i) deforming network, which starts from a regular 2D grid and adaptively deforms it into candidate UV coordinates; (ii) wrapping network, which maps those 2D UV candidates to the 3D surface by smoothly bending the deformed grid to conform to the input mesh, bridging the 2D and 3D domains; (iii) cutting network, which operates on the wrapped 3D surface to open the closed geometry by placing seams that reduce distortion for subsequent flattening; and (iv) unwrapping network, which projects the resulting open 3D surface back to 2D while preserving smoothness and geometric coherence. In this design, Each subnetwork is a pointwise MLP applied to mesh vertices, and the whole architecture is trained end-to-end in an unsupervised fashion. The four subnetworks form two complementary branches enforcing 2D–3D–2D and 3D–2D–3D cycles for ensuring consistency across 2D and 3D domains. Coupled with differentiable losses that encourage bijectivity and minimize distortion, this backbone produces candidate UV parameterizations for the next stages. Further details on the loss design and the interaction of the four sub-networks within the two-branch bi-directional cycle are given in Appendix A.2.
@@ 3.4 Visibility-Aware UV Parameterization
Next, we define a threshold $\tau$ based on the UV domain size. Specifically, we compute the bounding square of the UV map with side length $L(Q)$, and set $\tau=\tau_{\text{scale}}\cdot L(Q)$ where $\tau_{\text{scale}}$ is a constant (e.g., $0.1$ by default). The intuition behind Eq. 5 is that, during training, at each iteration we must identify the cutting-seam (boundary) points in UV space in order to evaluate their ambient occlusion values and then compute the visibility-aware loss in Eq. 7. Minimizing this loss encourages the model to learn UV parameterizations whose seams lie in less-visible (low-AO) regions. Eq. 5 provides a mechanism for detecting these seam points. Rather than locating seams directly on the 3D mesh, we detect them in the UV domain based on how local neighborhoods deform during unwrapping. Specifically, for each 3D vertex we know both its 3D coordinates and its corresponding UV coordinates at the current iteration. Therefore, we examine the 3D neighbors of each vertex and compare their mapped UV positions. If a pair of vertices are close neighbors on the 3D surface but become separated by more than a threshold distance in UV space, this indicates that the surface has been cut between them. We therefore mark such UV points as seam (boundary) points. This criterion is grounded in the fact that any valid unwrapping must introduce discontinuities (i.e., seams): vertices that are adjacent on the 3D surface become far apart in the UV plane after the mesh is cut open. Eq. 5 formalizes this idea by detecting precisely those locations where this large separation occurs. However, this maximization problem in Eq. 5 is not inherently differentiable, so we introduce mathematical modifications to ensure differentiability during the learning process. Specifically, instead of a hard decision, we assign each vertex a differentiable soft seam membership score: $s_{i}=\sigma\big(\beta\,(\eta_{i}-\tau)\big)$ where $\sigma(\cdot)$ is the sigmoid function and $\beta$ controls the sharpness of the transition. The maximization problem can then be formulated as a differentiable function:
@@
Visibility-Aware Seam Loss. After identifying soft seam memberships $s_{i}\in[0,1]$ for each vertex $v_{i}$, we encourage seams to lie in less visible regions of the surface. As a visibility proxy, we use per-vertex ambient occlusion values $AO_{i}\in[0,1]$, computed in the first step, where 0 indicates fully occluded and 11 indicates fully visible. The loss is defined as the weighted average of occlusion values, using soft seam memberships as weights:
@@
where $\varepsilon$ is a small constant to avoid division by zero. Intuitively, if the network assigns high seam weight $s_{i}$ to highly visible vertices (high $AO_{i}$), the loss is large. Conversely, assigning seam weight to occluded vertices (low $AO_{i}$) reduces the loss. Therefore, minimizing $\mathcal{L}_{\text{AO}}$ guides seam placement toward regions of low exposure.
@@ 4 Experiments and Results
We evaluate our unsupervised semantic- and visibility-aware UV parameterization pipelines on a diverse collection of meshes and compare against several baselines. We provide both qualitative and quantitative comparisons to: FlexPara Zhao et al. (2025) (recent state-of-the-art global parameterization network), Autodesk Maya Autodesk, Inc. (2025) (industry-grade commercial software), Blender Blender Foundation (2025) (popular open-source modeling tool), and xatlas jpcy (2025) (widely-used C++ library for generating unique texture coordinates suitable for lightmap baking and texture painting).
@@
Figs. 1,  3,  4, and  5 show representative qualitative results, and Tables. 1 and  3 presents quantitative results of our semantic-aware and visibility-aware pipelines against baselines (additional results are provided in Appendix A.3). In the following, we detail each experiment and its corresponding results.
4.1 Qualitative Evaluation
@@
Comparative Evaluation of Semantic-Aware UV Parameterization. To evaluate the performance of the proposed semantic-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025), Autodesk Maya’s Automatic UV generator Autodesk, Inc. (2025), Blender’s SmartUV Blender Foundation (2025), and xatlas jpcy (2025).Fig. 3 shows a representative comparison on a single Bunny mesh; additional results on more diverse and complex meshes are provided in Figs. 12 and 13 in Appendix A.3. In Fig. 3 and Fig. 12, each row corresponds to one method. Semantic parts are color-coded, and columns show the rendered 3D mesh from multiple camera viewpoints. The rightmost column depicts the unwrapped UV coordinates, where UV islands are color-coded consistently with their corresponding semantic 3D parts in the rendered views. As shown, our method successfully identifies semantic parts of the input mesh, parameterizes them, and maps them into UV space with low distortion. This leads to cleaner, more interpretable charts that directly benefit downstream tasks such as texture painting and transfer. By contrast, the baseline methods fail to consistently capture and parameterize semantic parts of the 3D mesh.
@@
Comparative Evaluation of Visibility-Aware UV Parameterization. To evaluate the performance of the proposed visibility-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025) and OptCuts Li et al. (2018). Figs. 4 and  9 present representative comparisons on different meshes. Additional examples appear in Figs. 5,  10, and 11 in Appendix A.3. In Fig. 4 the top row visualizes per-vertex ambient occlusion (AO): lighter tones (e.g., yellow) indicate high exposure (more visible) while darker tones (e.g., purple) indicate low exposure (more occluded). A desirable parameterization therefore places cutting seams on darker, less-visible regions to reduce perceptible seam artifacts after texturing and rendering. The rows beneath show UV parameterizations produced by our method, FlexPara, and OptCuts (in that order). Qualitatively, our method allocates a larger fraction of seam geometry to low-exposure regions compared to the baselines. Table 1 (and Fig. 8 in Appendix A.3) quantitatively supports this observation: the report the mean AO values of seam vertices for each method, where lower scores indicate seams located in less-visible (more occluded) areas. As shown, our method consistently attains lower mean seam-AO, confirming that the visibility-aware objective successfully biases seam placement toward less-exposed surface regions. Also, Fig. 5 illustrates checkerboard texturing on different mesh objects using parameterizations from our method and from baselines. As shown, our method produces noticeably more seamless texture patterns when rendered from visible viewpoints, whereas FlexPara and OptCuts exhibit prominent cutting seams (highlighted in the zoomed-in insets of Fig. 5).
4.2 Quantitative Evaluation
User Study. To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games.
Table 2: To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user‐study with 45 expert participants performing 11 comparisons between textured 3D shapes and UV parameterizations produced by our method and baselines. We report the percentage of participant preferences for each method. Our proposed method is strongly preferred by the expert users over the baselines.
Method Expert User Preference Percentage Visibility-Awareness Evaluation OptCuts Li et al. (2018) 1.48 % FlexPara (single-Chart) Zhao et al. (2025) 4.81 % Ours - Visibility-Aware Param 93.70 % Semantic-Awareness Evaluation xatlas jpcy (2025) 2.66 % Blender Blender Foundation (2025) 5.33 % Autodesk Maya Autodesk, Inc. (2025) 8.44 % FlexPara (multi-Chart) Zhao et al. (2025) 3.33 % Ours - Semantic-Aware Param 80.22 %
Each participant completed 11 comparisons between textured 3D shapes and UV parameterizations produced by our method, FlexPara Zhao et al. (2025), OptCuts Li et al. (2018), Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), and xatlas jpcy (2025). For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Table 2 reports the percentage of expert participant preferences for each method. As the table shows, expert users, particularly UV, layout, and modeling artists, strongly prefer our proposed method over the baselines. Examples of questions in our user‐study questionnaire are shown in Fig. 16. Also, Table 10 in Appendix A.3 reports the percentage of general participant preferences for each method. As shown in Table 10, general users also strongly prefer our proposed method over the baselines.
@@ 4.2 Quantitative Evaluation
Semantic- and Visibility-Awareness Evaluation. We also quantitatively evaluate both the semantic-aware and visibility-aware components of our pipeline on a collection of representative meshes. For the semantic-aware pipeline we report (i) Hamming distance (lower is better), the fraction of vertices whose labels disagree after optimal label matching to the reference (we use SAMesh Tang et al. (2024) outputs as a ground-truth); and (ii) Rand Index (RI) (higher is better), a pairwise agreement measure that counts how often vertex pairs are grouped or separated identically between the two labelings. For the visibility-aware pipeline we report Mean seam AO (lower is better in our AO convention, where 0 = occluded and 1 = exposed), the average ambient-occlusion value over seam vertices, which indicates how well seams are placed in occluded regions. Tables 1 and  3 summarize these metrics and compare our method against the baselines described in Sec. 4. Further discussion of the results in these two tables are provided in Appendix A.3.
Geometry-Preservation Evaluation. In addition to task-specific metrics mentioned above, in Tables 1 and  3, we evaluate geometric quality for all methods using conformality (angle-preservation) and equiareality (area-preservation) (higher is better) to ensure the parameterizations remain faithful enough to the original surface. The proposed semantic-aware and visibility-aware UV parameterization frameworks produce an increase in standard distortion metrics including angle-distortion and area-distortion. This is a trade-off which is expected when optimizing multiple objectives during the learning process. Specifically, our optimization introduces explicit semantic-alignment and visibility-awareness terms that encourage semantically consistent charts and reduce conspicuous seams. However, existing geometry-preserving methods (e.g., Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), xatlas jpcy (2025), and FlexPara Zhao et al. (2025)) focus only on traditional distortion objectives and therefore score better on those numbers, but they do not provide the semantic- and visibility-related guarantees produced by our method.
@@
We note that these semantic and visibility properties are often more valuable in some downstream tasks such as mass production, semantic editing, and 3D asset reuse, where consistent UV semantics and low seam visibility significantly reduce manual effort. Importantly, the trade-off is controllable in our framework: the relative emphasis between distortion and semantic/visibility objectives can be controlled by tunable weights and hyperparameters between different loss functions in Equations (1) and (8), so users can shift toward lower distortion when that is preferred.
@@ A.3 Additional Experiments and Results
We show expanded results, ablation studies, and discussions from the main paper featuring more surface parameterization and more viewpoints of different 3D objects in Figs. 716, and Tables. 5-  9.
Ablation studies on different UV packing (aggregator) methods. To further improve our semantic-aware UV parameterization pipeline’s output, a simple enhancement is to post-process the UV charts by jointly optimizing two key UV metrics: texel density and UV space utilization. Higher texel density improves the quality of the projected 3D texture, and maintaining consistent texel density across charts ensures uniform visual fidelity. At the same time, efficient space utilization maximizes the use of available texture memory. Hence, to analyze the effect of more advanced UV-chart packing (aggregator) strategies (different from the one proposed in Sec. 3.3), we perform an ablation study and evaluate several UV-packing techniques from three widely used commercial tools including, Autodesk Maya Autodesk, Inc. (2025), Houdini Toronto-based SideFX (2025), and Blender Blender Foundation (2025). We use a 3D cow mesh shown in Fig. 7 to evaluate texel density and UV-space utilization of UV parameterizations packed using different UV packing strategies from the three commercial tools. Specifically, we apply our semantic-aware UV parameterization framework as follows: (i) the 3D cow is first partitioned by PartField Liu et al. (2025) into seven semantic 3D parts, (ii) then the geometry-preserving neural backbone is applied separately to each part to learn UV parameterizations for each 3D part, (iii) instead of using our default UV packing (presented in Sec. 3.3), we apply different UV packing strategies from the three commercial tools, and (iv) lastly, for each packing strategy, we compute two UV metrics, namely texel density and UV-space utilization, for the UV parameterizations resulting from the three different UV packing methods. Table 5 reports the resulting values for both metrics to illustrate how each tool performs on a 3D cow mesh example. Also, in Fig. 7, we qualitatively compare the three commercial tools, applying their automatic packing and layout features to the same unwrapped mesh of 3D cow.
Table 5: Quantitative ablations of different UV-packing strategies from three widely used commercial tools: Autodesk Maya, Blender, and Houdini. We use a 3D cow mesh shown in Fig. 7 to evaluate texel density and UV-space utilization of UV parameterizations packed using different UV packing strategies from the three commercial tools. Specifically, after applying our semantic-aware UV parameterization framework, we obtain UV parameterizations for each segmented 3D part, resulting in seven UV charts. Then, instead of using our default UV packing (presented in Sec. 3.3), we apply different UV packing strategies from the three commercial tools. Lastly, for each packing strategy, we compute two UV metrics, namely texel density and UV-space utilization, for the UV parameterizations resulting from the three different UV packing methods. Higher texel density improves the quality of the projected 3D texture, and maintaining consistent texel density across charts ensures uniform visual fidelity. At the same time, efficient space utilization maximizes the use of available texture memory.
Method Texel Density UV Space Utilization Chart1 Chart2 Chart3 Chart4 Chart5 Chart6 Chart7 Avg Maya Autodesk, Inc. (2025) 160.62 150.41 150.41 150.41 156.51 150.41 150.41 152.74 71% Blender Blender Foundation (2025) 151.66 142.02 142.02 142.02 147.78 142.02 142.02 144.22 57% Houdini Toronto-based SideFX (2025) 157.08 147.09 147.09 147.09 153.06 147.09 147.10 149.37 62%
Figure 7: Qualitative ablations of different UV-packing strategies. To further improve our semantic-aware UV parameterization pipeline’s output, we present alternative more advanced UV packing results (different from the one proposed in Sec. 3.3) obtained from three widely used commercial tools: (a) Autodesk Maya, (b) Blender, and (c) Houdini. The left-most column shows the 3D cow mesh segmented by PartField Liu et al. (2025). In the Autodesk Maya result (a), each UV chart corresponding to a semantic 3D part is colored to match its counterpart on the 3D mesh. The same applies to the Blender and Houdini results. Specifically, the 3D cow is first partitioned by PartField Liu et al. (2025) into seven semantic 3D parts. Then, the geometry-preserving neural backbone (Sec. 3.2) is applied separately to each part to learn UV parameterizations for each 3D part leading to seven UV charts. Then, instead of using our default UV packing (presented in Sec. 3.3), we apply automatic packing strategies layout features of the three commercial tools to the unwrapped mesh of 3D cow.
Ablation studies on different partitioning strategies (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize (semantic-aware UV parameterization) pipeline identical. Then, we train each pipeline separately on input 3D meshes and measure and report angle-preservation (conformality) and area-preservation (equiareality) properties in the generated UV parameterization. Figs. 1415 and Table 6 shows the qualitative and quantitative results of these experiments, respectively. To interpret the numbers in the table, we point out two points: (i) There is no obvious or general relationship between the choice of segmentation algorithm and the geometry-preservation metrics (conformality and equiareality) of the final UV parameterization. This is due to the fact that existing 3D segmentation methods were developed to produce semantically meaningful parts, not to optimize downstream UV parameterization. Consequently, they do not incorporate UV-specific information. Therefore, from the numbers in Table 6, we cannot draw a unified, causal conclusion connecting segmentation method type to final UV geometry properties. Having that said, we include this segmentation ablation and Table 6 intentionally to inform practitioners that different partitioning methods, while varying in semantic quality, can also influence geometry-preservation in the resulting UVs. Reporting these results shows the important trade-off between semantic-awareness and geometric fidelity. Therefore, practitioners should try different segmentation approaches and choose the one that best fits their application needs. (ii) Of course, state-of-the-art, learning-based segmentation methods can produce more semantically coherent partitions than ShDF Shapira et al. (2008) (as shown by works such as SAMesh Tang et al. (2024)). Therefore, any improvements in that literature can directly improve the semantic-awareness of our pipeline because segmentation is the first stage of our semantic-aware parameterization (Fig.2). In other words, a better 3D partitioner is an immediate plug-in improvement for our method.
Ablation studies on geometry-preservation vs. semantic / visibility objectives. Optimizing multiple geometry-preservation metrics simultaneously (conformality, equiareality, stretch, bijectivity, etc.) is known to require trade-offs. Specifically, achieving the best possible value for every metric at once is generally impractical. Introducing higher-level objectives (semantic- and visibility-awareness), that we propose in this work, increases this tension and therefore can further degrade some geometry metrics. To quantify these effects, we includ an ablation in Table. 7 that compares our proposed pipelines to the baseline (labelled “Base Neural UV Arch” in Table. 7, i.e., FlexPara) which is state-of-the-art for geometry preservation. We compares three variants of our pipeline to the baseline: (i) Semantic-Aware Param, the pipeline of Sec. 3.3 (Fig. 2), which first partitions the mesh into semantic parts and then learns per-part parameterizations; (ii) Visibility-Aware Param, the pipeline of Sec. 3.4 (Algorithm. 1), which produces a global atlas while steering seams toward less-visible regions; and (iii) Visibility+Semantic-Aware Param, the combined pipeline that applies both sets of objectives. Three observations explain the quantitative trends: (i) when comparing conformality and equiareality, the Semantic-Aware and the combined Visibility+Semantic-Aware pipelines lose little equiareality relative to the baseline and even show slight improvement in conformality in some cases. Intuitively, this is because the semantic-aware stage first partitions the mesh into multiple semantically meaningful subparts. Intuitively, learning a geometry-preserving parameterization on smaller, simpler patches is easier. In other words, the network can better satisfy geometry-preservation while also enforcing semantic consistency. Therefore, semantic partitioning provides a natural way to reduce distortion without heavily sacrificing geometry metrics. (ii) By contrast, the visibility-aware pipeline does not partition the mesh. Instead, it steers seams to less-visible (more occluded) regions while producing a global atlas (a single large chart). Unwrapping a whole 3D mesh into a single UV chart to while simultaneously optimizing visibility and geometry is significantly harder. Consequently, the Visibility-Aware pipeline shows a larger drop in equiareality compared to the baseline and compared to the semantic-aware pipelines. (iii) The combined Visibility+Semantic-Aware pipeline aims to obtain the best of both worlds: it first partitions the mesh into smaller parts to retain geometry preservation, and then steers seams toward less-visible regions within those parts. This combined approach yields a more favorable trade-off: it loses far less equiareality (0.6369) than the pure visibility-aware pipeline (0.6093) while still improving seam visibility and semantic consistency.
Analysis on training and inference computation times. To provide users and practitioners with some insights about the computation time needed for the proposed method in this study, we report both inference and training time of the proposed method in Table 3, Table 1, Table 8, and Table 9, and compare them with the computation time of the state-of-the-art method (at the time of writing this manuscript), FlexPara Zhao et al. (2025).
  • Inference time. (i) Visibility-aware pipeline: although training is performed per-mesh (as is common in this literature: Nuvo Srinivasan et al. (2024), FlexPara Zhao et al. (2025), OptCuts Li et al. (2018)), the visibility objective is applied only during training (not the inference stage). Specifically, at inference the trained model uses the same neural architecture as FlexPara Zhao et al. (2025) and therefore incurs no additional runtime overhead. This is why the inference times for our Visibility-Aware Param and FlexPara are the same in Table 1 ($\approx$2 seconds). (ii) Semantic-aware pipeline: the semantic-aware pipeline does introduce extra inference cost because it performs three stages at test time: (a) 3D semantic partitioning, (b) per-part parameterization (running the backbone once per part), and (c) aggregation/UV packing. These additional steps (segmentation, multiple backbone invocations, and packing) explain the larger inference time reported in Table 2 and for the combined pipeline (semantic+visibility-aware) in Table 1. A simple and effective mitigation is to parallelize the per-part backbone calls by using multi-threading during the inference, which would significantly reduce inference time.
  • Training time. (i) Semantic-aware pipeline (Table 8): counterintuitively, our semantic-aware pipeline achieves shorter training time than FlexPara Zhao et al. (2025) across different 3D meshes. The reason is that the mesh is partitioned into many smaller submeshes: although our semantic-aware pipeline calls the backbone multiple times (once per semantic part), which might suggest longer training, this is compensated by the fact that the 3D segmented parts are much smaller than the original input mesh used by FlexPara. Hence, training on many small submeshes is faster than training a single model on a large, complex mesh with many vertices and faces. This empirical observation is consistent with our careful time measurement of each stage in the semantic-aware pipeline which is: learning UV parameterization for small 3D parts is much faster than doing the same for large meshes. More importantly, the reduction in per-part training cost outweighs the multiplicative overhead of multiple backbone runs, which results in shorter overall training time for the semantic-aware pipeline in our experiments. This is why Table 8 shows shorter training times for our semantic-aware pipeline compared to FlexPara in our experiments. (ii) Visibility-aware pipeline (Table 9): The visibility-aware pipeline trains slower than FlexPara Zhao et al. (2025) across different meshes. This is expected because, compared to the base neural backbone (i.e., FlexPara), our visibility-aware pipeline introduces per-iteration overheads: (a) per-vertex ambient occlusion (AO) computation, (b) candidate UV generation via the backbone, (c) seam/boundary extraction in UV space, and (d) AO-weighted loss computation. The most expensive stage is seam extraction which accounts for approximately 90% of the additional training time. Specifically, for each vertex we inspect its 3–5 neighbors, compare mapped UV positions, and evaluate Eq. 5 to detect UV discontinuities. This neighbor-checking and distance computation scales with vertex count and dominates the additional training cost ( 90% of the overhead), making visibility-aware training significantly heavier for complex meshes.
  • Potential attempts to overcome computation burdens. To mitigate the computational burdens discussed above, we propose two practical solutions: (i) parallelizing per-part backbone calls in semantic-aware pipeline. In our current implementation, the neural backbone is executed sequentially for each semantic part. Switching to a multi-threaded or multi-GPU implementation to run per-part backbone calls in parallel would significantly reduce both inference and training times. (ii) Optimizing seam detection in visibility-aware pipeline. The seam-detection stage is the main training-time bottleneck. Algorithmic and engineering improvements, e.g., spatial hashing, cached neighborhood lookups, and GPU-accelerated neighbor/distance checks, can significantly reduce this overhead. We note these optimizations as straightforward improvements and sensible directions for future work.
@@
Table 6: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Quantitatively, using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh.